OpenStudy (anonymous):

Simplify: 9a^2/3a+b - b^2/3a+b

OpenStudy (raden):

the denominator was same, what should u do, first ?

OpenStudy (anonymous):

cross out

OpenStudy (raden):

so, what do u get ?

OpenStudy (anonymous):

9a^2-b^2 ?

OpenStudy (raden):

now, can u factor it ?

OpenStudy (anonymous):

???

OpenStudy (raden):

u must factorise the term of 9a^2-b^2 hint : a^2 - b^2 = (a+b)(a-b)

OpenStudy (anonymous):

oops I mean (9+b)(9-b^2)

OpenStudy (raden):

nopes.... it should be 9a^2-b^2 = (3a+b)(3a-b)

OpenStudy (anonymous):

the 3 is from the denominator?

OpenStudy (raden):

not always... 9a^2-b^2 = (3a)^3 - b^2 = (3a+b)(3a-b) so, what should u do if the expression becomes